OpenStudy (anonymous):

solve for k : p=2k+30

OpenStudy (anonymous):

brother dont you think u shud give two equations?

OpenStudy (anonymous):

2k is 2times k + 30

OpenStudy (anonymous):

p = 2k + 30 -- subtract 30 from both sides p - 30 = 2k -- now divide both sides by 2 (p - 30)/2 = k 1/2p - 15 = k